Explain briefly how each of the following transactions would affect a company’s balance sheet. Remember, assets must equal liabilities plus owners’ equity before and after the transaction.

a) Sale of used equipment with a book value of $300,000 for $500,000 cash.

b) Purchase of a new $80 million building, financed 40 percent with cash and 60 percent with a bank loan.

c) Purchase of a new building for $60 million cash.

d) A $40,000 payment to trade creditors.

e) A firm’s repurchase of 10,000 shares of its own stock at a price of $24 per share.

f) Sale of merchandise for $80,000 in cash.

g) Sale of merchandise for $120,000 on credit.

h) Dividend payment to shareholders of $50,000
Answer #1

(a)When equipment is sold for 500000 with book value 300000 then fixed assets will be decreased by 300000 and cash will be increased by 500000 and owners equity will be increased by profit which is 200000

Hence if we talk about net effect then assets increased by 500000-300000*=200000

And liabilities and equity increased by 200000

So before and after transaction ,assets = liabilities+equity

b.In this case,fixed assets increased by 80 million and cash or bank decreased by 80*0.4=32 million and liabilities increased by 48 million because of increase in bank loan

Hence net effect on assets side =80-32=48 million increment

And effect on liabilities and equity side also 48 million

increment

C.In this case ,fixed assets will be increased by 60 million because building purchased and at the same time ,cash will be decreased by 60 million because building purchased in cash

Hence net effect on assets side =60-60=0

And effect on liabilities and equity side is also 0

d.Whem payment made to trade creditors then cash will be decreased by 40000 and at the same time trade creditors will be decreased by 40000.

Therefore ,assets side and liabilities & equity side both decreased by 40000